### v3

# When are off-diagonal hypergraph Ramsey numbers polynomial?

Abstract
A natural open problem in Ramsey theory is to determine those $3$-graphs $H$ for which the off-diagonal Ramsey number $r(H,K_{n}^{(3)})$ grows polynomially with $n$. We make substantial progress on this question by showing that if $H$ is tightly connected or has at most two tight components, then $r(H,K_{n}^{(3)})$ grows polynomially if and only if $H$ is contained in an iterated blowup of an edge.
1 Introduction
Given a $k$-uniform hypergraph $H$ (henceforth, $k$-graph), the off-diagonal Ramsey number $r(H,K_{n}^{(k)})$ is the smallest natural number $N$ such that every red/blue-coloring of the edges of $K_{N}^{(k)}$, the complete $k$-graph with $N$ vertices, contains either a red copy of $H$ or a blue copy of $K_{n}^{(k)}$. For graphs, the $k=2$ case, we know that $r(H,K_{n})$ always grows polynomially with $n$ and the main problem is to determine the growth rate more precisely. This problem remains open even when $H$ is a clique, where the correct polynomial dependency is only understood for $K_{3}$ and $K_{4}$ — for $K_{3}$ the off-diagonal Ramsey number was famously determined up to a constant factor by Ajtai, Komlós and Szemerédi [1] and Kim [12], while for $K_{4}$ a recent result of Mattheus and Verstraëte [13] shows that $r(K_{4},K_{n})=n^{3+o(1)}$. When $H$ is a cycle, even less is known and determining the polynomial order of $r(C_{4},K_{n})$ in $n$ is a major Erdős problem (see [3, 4] for the best bounds on this problem and [9, 16] for recent progress on other cycle-complete Ramsey numbers).
For $3$-graphs $H$, $r(H,K_{n}^{(3)})$ does not always grow polynomially. Indeed, it is already the case [8] that $r(K_{4}^{(3)},K_{n}^{(3)})\geq 2^{\Omega(n\log n)}$. Much of the recent work on off-diagonal hypergraph Ramsey numbers has focused on extending this lower bound to other hypergraphs, so that it is now known that $r(H,K_{n}^{(3)})\geq 2^{\Omega(n\log n)}$ for $K_{4}^{(3)}\setminus e$ and, more generally, all links of odd cycles [11] and for all tight cycles of length not divisible by three [6].
Here we look in a different direction, our concern being with the problem of classifying those $H$ for which $r(H,K_{n}^{(3)})$ grows polynomially in $n$. Although this problem has been studied since at least work of Erdős and Hajnal [10] in the early 1970s, it seems to have been first raised explicitly by the first author [5] at an AIM workshop in 2015. We propose a full classification, as follows.
Conjecture 1.1.
For a $3$-graph $H$, there exists a constant $c$ depending only on $H$ such that $r(H,K_{n}^{(3)})\leq n^{c}$ for all $n$ if and only if $H$ is a subgraph of an iterated blowup of an edge.
To clarify what we mean by an iterated blowup, we first note that a blowup of an edge is simply a complete tripartite $3$-graph. An iterated blowup of an edge is then any graph which is either a blowup of an edge or formed iteratively by placing another iterated blowup in one or more of the parts in a blowup of an edge. In what follows, we say that a $3$-graph is iterated tripartite if it is contained in an iterated blowup of an edge.
The direction of Conjecture 1.1 saying that there exists a constant $c$ such that $r(H,K_{n}^{(3)})\leq n^{c}$ if $H$ is iterated tripartite was already known to Erdős and Hajnal [10] (see also [8, 11]). Their concern was with a slightly different Ramsey-type question. For natural numbers $4\leq s<n$ and $2\leq t\leq\binom{s}{3}$, they were interested in the Ramsey function $r_{3}(s,t;n)$, the smallest natural number $N$ such that every red/blue-coloring of the edges of $K_{N}^{(3)}$ contains either a blue copy of $K_{n}^{(3)}$ or a set of $s$ vertices with at least $t$ red edges. Regarding the behavior of this function, they conjectured that there should be a polynomial-to-exponential transition for the growth rate of $r_{3}(s,t;n)$ at $t=t(s)$, the maximum number of edges in an iterated tripartite $3$-graph with $s$ vertices. Their conjecture may be seen as a toy model for our Conjecture 1.1.
In their paper, Erdős and Hajnal [10] proved one direction of their conjecture, showing that $r_{3}(s,t;n)$ grows polynomially in $n$ for $t\leq t(s)$. Their proof, as indicated above, further shows one direction of our Conjecture 1.1. However, it remains open to show that $r_{3}(s,t;n)$ grows exponentially in a power of $n$ for $t>t(s)$. Some partial results, saying, for instance, that the conjecture holds when $s$ is a power of $3$, were proven by Conlon, Fox and Sudakov [8], while the analogous problem in higher uniformities was solved completely by Mubayi and Razborov [15]. In particular, we see that Conjecture 1.1 holds if $|H|$ is a power of $3$ and the number of edges in $H$ is larger than that in any iterated blowup with $|H|$ vertices.
In light of these results, it is perhaps surprising that Conjecture 1.1 has not been stated before. One reason for this was the common belief in the community that $r(H,K_{n}^{(3)})$ should also be polynomial for linear $H$, where a hypergraph is said to be linear if any two edges in the hypergraph share at most one vertex. However, it was recently shown [6] that this belief is mistaken and that there are linear hypergraphs for which $r(H,K_{n}^{(3)})$ grows superpolynomially. Conjecture 1.1 is stronger again, suggesting that linearity is a red herring in this context.
Towards Conjecture 1.1, we prove two results. We say that a $3$-graph $H$ is tightly connected if, for any two edges $e$ and $f$ of $H$ there exists a sequence of edges $e=e_{0},e_{1},\dots,e_{t}=f$ such that $e_{i-1}$ and $e_{i}$ share two vertices for all $i=1,\dots,t$. Our first result says that Conjecture 1.1 holds for tightly connected hypergraphs, even giving a lower bound in this case which is exponential in a power of $n$.
Theorem 1.2.
If $H$ is a $3$-graph which is tightly connected and not tripartite, then $r(H,K_{n}^{(3)})\geq 2^{\Omega(n^{2/3})}$.
One objection towards this being strong evidence for Conjecture 1.1 is that a tightly connected hypergraph is almost exactly the opposite of a linear hypergraph. Our second result goes some small way towards overruling this objection. We define a tight component of $H$ to be a maximal tightly connected subgraph and observe that the edge set of every $3$-graph decomposes into disjoint tight components. We prove that Conjecture 1.1 holds for hypergraphs which have at most two tight components, though our lower bound is considerably weaker in this case.
Theorem 1.3.
If $H$ is a $3$-graph with at most two tight components and not iterated tripartite, then $r(H,K_{n}^{(3)})\geq 2^{\Omega(\log^{2}n)}$.
The proof of Theorem 1.3 is significantly more technical than that of Theorem 1.2 and may be considered our main result. With that in mind, we will warm up by proving Theorem 1.2 in the next section, before returning to Theorem 1.3 in Section 3. We then conclude with some further remarks and open problems.
2 Tightly connected hypergraphs
In this short section, we prove Theorem 1.2, the statement that if $H$ is tightly connected and not tripartite, then $r(H,K_{n}^{(3)})\geq 2^{\Omega(n^{2/3})}$. In fact, we will prove a stronger statement, from which 1.2 clearly follows.
Theorem 2.1.
For every positive integer $N$, there is a red/blue edge coloring of $K_{N}^{(3)}$ vertices such that any red tightly connected subgraph is tripartite and the largest blue clique has order $O((\log N)^{3/2})$.
Proof of Theorem 2.1.
Let $\ell=C\log N$, where $C>0$ is a sufficiently large absolute constant, and $n=\ell^{3/2}$. Let $r=0.01\ell$ and let $V$ be an $r$-trifference code in $\{1,2,3\}^{\ell}$ of size $N$. What this means is that $V$ is a subset of $\{1,2,3\}^{\ell}$ of size $N$ with the property that, for any triple uvwuvw of elements of $V$, there are at least $r$ coordinates where $\{u_{i},v_{i},w_{i}\}=\{1,2,3\}$. Such a set of size $N$ can be found by the first moment method, as it is exponentially unlikely that a random triple uvwuvw is not $r$-trifferent.
We will define our Ramsey coloring on the complete $3$-graph with vertex set $V$. Let $c(uv)$ denote the set of coordinates in which $u,v$ differ. By the definition of $V$, $r\leq|c(uv)|\leq\ell$ for all $u,v$. Define $\phi(uv)$ to be an element of $c(uv)$ picked uniformly at random. The Ramsey coloring of $K_{N}^{(3)}$ will be $\chi$ where $\chi(uvw)$ is red if and only if $\phi(uv)=\phi(vw)=\phi(uw)$. Observe that, by definition, any red triple uvwuvw automatically satisfies $\{u_{i},v_{i},w_{i}\}=\{1,2,3\}$ for $i=\phi(uv)$.
Red tight components are tripartite. For any red tight component, all pairs in its $2$-shadow must share the same $\phi$-value $i$. Then the $i$-th coordinates of the vertices give a tripartition of the vertices, showing that the component is tripartite.
No blue $K_{n}^{(3)}$ with positive probability. Fix a set $U$ of $n$ vertices in $V$; we would like to bound the probability that this $n$-set forms a blue clique. Let $T_{uvw}$ be the event that the triple uvwuvw is red, that is, that the triangle uvwuvw is monochromatic under $\phi$. Each event has probability
$\Pr[T_{uvw}]=\frac{|c(uv)\cap c(vw)\cap c(uw)|}{|c(uv)|\cdot|c(vw)|\cdot|c(uw)|}=\Theta(\ell^{-2})$
by our choice of $V$ and $c(\cdot)$. Since there are $\Theta(n^{3})$ total events, the expected number of monochromatic triangles in $U$ is $\mu=\Theta(n^{3}/\ell^{2})=\Theta(\ell^{5/2})$. The Poisson paradigm predicts that
$\Pr[\bigwedge\overline{T}_{uvw}]\leq e^{-\Theta(\ell^{5/2})}.$ (1)
We first show that this would suffice to prove the theorem. Indeed, since there are at most $N^{n}$ $n$-sets and
$N^{n}e^{-\Theta(\ell^{5/2})}<1$
by our choices of $n,\ell$ and $N$, the union bound gives that there is a positive probability no blue $n$-cliques appear.
Now we prove (1) via Suen’s inequality in the form given in Alon–Spencer [2, Theorem 8.7.1], which implies that
$\Pr[\bigwedge\overline{T_{\tau}}]\leq e^{-\mu+\sum_{\tau\sim\tau^{\prime}}y(\tau,\tau^{\prime})},$ (2)
where $\tau$ ranges over all triangles uvwuvw in $U$ and $\tau\sim\tau^{\prime}$ if these two triangles share at least one edge, with
$y(\tau,\tau^{\prime})=(\Pr[T_{\tau}\wedge T_{\tau^{\prime}}]+\Pr[T_{\tau}]\cdot\Pr[T_{\tau^{\prime}}])\prod_{\tau^{\prime\prime}\sim\tau\textnormal{ or }\tau^{\prime\prime}\sim\tau^{\prime}}(1-\Pr[T_{\tau^{\prime\prime}}])^{-1}.$
Observe that if $\tau$ and $\tau^{\prime}$ share an edge, then $\Pr[T_{\tau}\wedge T_{\tau^{\prime}}]=O(\ell^{-4})$ and $\Pr[T_{\tau}]\cdot\Pr[T_{\tau^{\prime}}]=O(\ell^{-4})$. Moreover, there are $\Theta(n)$ triangles $\tau^{\prime\prime}$ that share an edge with either $\tau$ or $\tau^{\prime}$, so we obtain
$y(\tau,\tau^{\prime})=O(\ell^{-4})(1-\Theta(\ell^{-2}))^{-\Theta(n)}=O(\ell^{-4})e^{\Theta(n/\ell^{2})}=O(\ell^{-4}),$
where we used that $n=o(\ell^{2})$. Since there are $O(n^{4})$ pairs of triangles $\tau,\tau^{\prime}$ that share an edge, (2) reduces to
$\Pr[\bigwedge\overline{T_{\tau}}]\leq e^{-\mu+O(n^{4}/\ell^{4})}=e^{-\Theta(\ell^{5/2})+O(\ell^{2})}=e^{-\Theta(\ell^{5/2})},$
as desired. ∎
Before moving on to the case with two tight components, we sketch an alternative, but quantitatively weaker, construction for the single component case. Let $\ell=C\log N$ and $n=C^{7/4}\log^{2}N$ for $C$ a sufficiently large constant. For each edge $uv$ of the complete graph on $V=[N]$, label it with some $\phi(uv)\in\{1,2,$$\ell\}$ chosen independently and uniformly at random. Moreover, assign a string $f(v)\in\{1,2,3\}^{\ell}$ to each vertex $v$, again chosen independently and uniformly at random. The Ramsey coloring of $K_{N}^{(3)}$ will be $\chi$ where $\chi(uvw)$ is red if and only if
  1. 1.
    $\phi(uv)=\phi(vw)=\phi(uw)=i$ for some $i$,
  2. 2.
    $f_{i}(u),f_{i}(v)$ and $f_{i}(w)$ are all distinct for this $i$
and blue otherwise. If there is any red tightly connected $3$-graph in $\chi$, then all edges in its $2$-shadow have the same $\phi$-value $i$, so the $f_{i}(v)$ yield a tripartition. To show that there is no blue $K_{n}$ with positive probability, we use an application of Janson’s inequality to show that, with positive probability, for every subset $U$ of $V$ with $n$ vertices and every $i\in\{1,2,$$\ell\}$ there are at least $n/6$ vertex-disjoint triangles in $U$ which are monochromatic of color $i$ under $\phi$. If we now bring the randomness of $f$ into play, then each such monochromatic triangle forms a blue edge with probability $7/9$. As the randomness for each coordinate $f_{i}$ of $f$ is independent, the induced hypergraph on any given $n$-vertex set $U$ is completely blue with probability at most $(7/9)^{\frac{n}{6}\cdot\ell}=\exp(-\Theta(n\ell))$. Since there are ${N\choose n}=\exp(O(n\log N))$ subsets to consider, the union bound and the fact that $\ell$ dominates $\log N$ completes the analysis, yielding the bound $r(H,K_{n}^{(3)})\geq 2^{\Omega(n^{1/2})}$ for every non-tripartite tightly connected $3$-graph $H$.
3 Two tight components
In this section, we prove Theorem 1.3, which says that if $H$ has at most two tight components and is not iterated tripartite, then $r(H,K_{n}^{(3)})\geq 2^{\Omega(\log^{2}n)}$. We will once again prove 1.3 by providing a single construction avoiding all such $H$ in red simultaneously.
Theorem 3.1.
For every positive integer $N$, there is a red/blue edge coloring of $K_{N}^{(3)}$ such that any red subgraph with at most two tight components is iterated tripartite and the largest blue clique has order $\exp(O(\sqrt{\log N}))$.
For the proof, we will need to show the existence of an edge coloring of the complete graph with certain powerful properties, encapsulated in the following lemma.
Lemma 3.2.
For positive integers $A\geq 20$, $\ell\geq 1$ and $N=2^{\ell}$, there exists an edge coloring $\varphi:E(K_{N})\rightarrow[\ell A]$ such that:
  1. 1.
    For each color, the edges with that color form a vertex-disjoint union of bicliques.
  2. 2.
    In every $k$-vertex subset with $k=\exp\left(\Omega(\log N/\log A)\right)$, there are $\Omega(k^{2.5})$ rainbow triangles.
We will postpone the proof of this lemma for now, first showing how to use it to prove Theorem 1.3.
Proof of Theorem 3.1.
Replace $N$ by the smallest power of $2$ exceeding $N$ and let $\ell=\log_{2}N$. Let $\varphi$ be as in Lemma 3.2 with $A,k=\exp(\Theta(\sqrt{\ell}))$. Let $g,f_{1},f_{2},f_{3}$ be four independent random functions on $N]\choose 2$}, where $g:{[N]\choose 2}\rightarrow\{1,2,3\}$ uniformly at random and $f_{1},f_{2},f_{3}:{[N]\choose 2}\rightarrow[\ell A]$ uniformly at random. The Ramsey coloring of the complete $3$-graph on $[N]$ will be $\chi$ where $\chi(uvw)$ for a triple uvwuvw with $u<v<w$ is red if and only if
  1. 1.
    uvwuvw is a rainbow triangle with respect to $\varphi$,
  2. 2.
    $g(uv)=1$, $g(vw)=2$ and $g(uw)=3$,
  3. 3.
    $f_{1}(uw)=f_{1}(vw)=\varphi(uv)$,
  4. 4.
    $f_{2}(uv)=f_{2}(uw)=\varphi(vw)$,
  5. 5.
    $f_{3}(uv)=f_{3}(vw)=\varphi(uw)$.
Otherwise, the triple uvwuvw is colored blue.
No large blue cliques.
We upper bound the probability that some subset $S$ of order $k$ forms a blue clique. Provided $k=\exp(\Omega(\ell/\log A))$, by the properties of $\varphi$, there are $\Omega(k^{2.5})$ rainbow triangles in $S$. By greedily picking out triangles, we may find $\Theta(k^{1.5})$ of them that are edge-disjoint. Each such triangle is colored red independently with probability $\Theta((\ell A)^{-6})$. Therefore, the probability that none of them is colored red is $e^{-\Theta(k^{1.5}(\ell A)^{-6})}.$ By the union bound, the probability that there is a blue clique can then be bounded by
$\binom{N}{k}e^{-\Theta(k^{1.5}(\ell A)^{-6})}=e^{\Theta(k\log N)-\Theta(k^{1.5}(\ell A)^{-6})},$
which is less than $1$ if $k=\Omega(\ell^{14}A^{12})$. To balance this with the condition that $k=\exp(\Omega(\ell/\log A))$, it suffices to pick $A,k=\exp(\Theta(\sqrt{\ell}))$, as indicated at the outset.
Union of two red tight components is iterated tripartite.
We now show that the union of any two red tight components is iterated tripartite. We start by proving a structural result for each red tight component.
Lemma 3.3.
Every red tight component $\mathcal{C}(V_{\mathcal{C}},E_{\mathcal{C}})$ is tripartite and there is a unique tripartition $V^{(1)}_{\mathcal{C}}\cup V^{(2)}_{\mathcal{C}}\cup V^{(3)}_{\mathcal{C}}$ of the vertices $V_{\mathcal{C}}$. Moreover, there exist three colors $c^{(1)}_{\mathcal{C}}$, $c^{(2)}_{\mathcal{C}}$ and $c^{(3)}_{\mathcal{C}}$ such that, for any ordering $i,j,k$ of $1,2,3$, $\varphi(uv)=c^{(k)}_{\mathcal{C}}$ for any $u\in V^{(i)}_{\mathcal{C}}$ and $v\in V^{(j)}_{\mathcal{C}}$.
Proof.
For any hyperedge $\{x,y,z\}$, define $\Phi(\{x,y,z\})$ to be the set $\{\varphi(xy),\varphi(yz),\varphi(xz)\}$. For any two red hyperedges $e,e^{\prime}$ sharing two vertices $x,y$, if $e=\{x,y,z\}$ and $e^{\prime}=\{x,y,z^{\prime}\}$, then the relative order of $x,y,z$ and the relative order of $x,y,z^{\prime}$ have to be the same, showing that $g(xz^{\prime})=g(xz)$ and $g(yz^{\prime})=g(yz)$. Therefore,
$\varphi(xz)=f_{g(xz)}(xy)=f_{g(xz^{\prime})}(xy)=\varphi(xz^{\prime})$
and similarly $\varphi(yz)=\varphi(yz^{\prime})$. Hence, as $\mathcal{C}$ is tightly connected in red, $\Phi(e)$ is the same for any $e\in E_{\mathcal{C}}$.
Now pick an arbitrary $e\in E_{\mathcal{C}}$ and let $c^{(1)}_{\mathcal{C}},c^{(2)}_{\mathcal{C}},c^{(3)}_{\mathcal{C}}$ be an arbitrary enumeration of the elements in $\Phi(e)$. Moreover, let $G$ be the subgraph of $K_{N}$ obtained by taking the $2$-shadow of $\mathcal{C}$. By induction, it is easy to see that the edges of color $c^{(i)}_{\mathcal{C}}$ in $G$ are connected for each $i\in[3]$.
Now, for any $v\in V_{\mathcal{C}}$, choose arbitrary vertices $u,w$ such that $\{u,v,w\}\in E_{\mathcal{C}}$ and define $i(v)$ to be such that $\varphi(uw)=c^{i(v)}_{\mathcal{C}}$. We show that $i(v)$ does not depend on the choice of $u$ and $w$. Suppose instead that there are some other $u^{\prime},w^{\prime}$ with $\{u^{\prime},v,w^{\prime}\}\in E_{\mathcal{C}}$ such that $\varphi(uw)\neq\varphi(u^{\prime}w^{\prime})$. Then we know that one of $\varphi(u^{\prime}v)$ and $\varphi(vw^{\prime})$ is equal to $\varphi(uw)$. Thus, in $G$, $u,v,w$ all belong to the connected subgraph of color $\varphi(uw)$ with $u$ and $w$ neighbors. Therefore, without loss of generality, we can assume that there is a monochromatic odd walk from $v$ to $u$ of color $\varphi(uw)$. Since the edges with color $\varphi(uw)$ form a vertex-disjoint union of bicliques, it must then be the case that $\varphi(uv)=\varphi(uw)$. But this is a contradiction, so $i(v)$ must be independent of the choice of $u$ and $w$.
We partition $V_{\mathcal{C}}$ into $V^{(1)}_{\mathcal{C}},V^{(2)}_{\mathcal{C}}$ and $V^{(3)}_{\mathcal{C}}$ based on the labels $i(v)$. It is clear that $\mathcal{C}$ is tripartite with respect to this tripartition and, as $\mathcal{C}$ is tightly connected, that this is the only possible tripartition. Finally, if $i,j,k$ is an ordering of $1,2,3$, then it is also clear that the connected subgraph of $G$ of color $c^{(k)}_{\mathcal{C}}$ is a bipartite graph on $V^{(i)}_{\mathcal{C}}\cup V^{(j)}_{\mathcal{C}}$. If $u\in V^{(i)}_{\mathcal{C}}$ and $v\in V^{(j)}_{\mathcal{C}}$, then, by the fact that the edges of color $c^{(k)}_{\mathcal{C}}$ form a vertex-disjoint union of bicliques in $K_{N}$, we know that $\varphi(uv)=c^{(k)}_{\mathcal{C}}$, as desired. ∎
Suppose now that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are two red tight components whose union is not tripartite. Without loss of generality, this means that there is some $u\in V^{(1)}_{\mathcal{C}}\cap V^{(1)}_{\mathcal{C}^{\prime}}$ and $v\in V^{(2)}_{\mathcal{C}}\cap V^{(1)}_{\mathcal{C}^{\prime}}$. If some $w\in V^{(3)}_{\mathcal{C}}$ is in $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$, then, since every edge between $V^{(1)}_{\mathcal{C}^{\prime}}$ and $V^{(i)}_{\mathcal{C}^{\prime}}$ has the same color under $\varphi$ for each of $i=2,3$, $c^{(2)}_{\mathcal{C}}=\varphi(uw)=\varphi(vw)=c^{(1)}_{\mathcal{C}}$, which is a contradiction. Therefore, $V^{(3)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must be empty. If there is some $w\in V^{(1)}_{\mathcal{C}}$ that is in $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$, say $w\in V^{(2)}_{\mathcal{C}^{\prime}}$, then, since every edge between $V^{(1)}_{\mathcal{C}}$ and $V^{(2)}_{\mathcal{C}}$ has the same color under $\varphi$, $\varphi(uv)=c^{(3)}_{\mathcal{C}}=\varphi(vw)=c^{(3)}_{\mathcal{C}^{\prime}}$. But then picking any $x\in V^{(2)}_{\mathcal{C}^{\prime}}$ we have a monochromatic triangle uvxuvx, which contradicts the fact that each color class in $\varphi$ is bipartite. Therefore, $V^{(1)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must also be empty. Similarly, $V^{(2)}_{\mathcal{C}}\cap(V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}})$ must be empty. Hence, $V^{(2)}_{\mathcal{C}^{\prime}}\cup V^{(3)}_{\mathcal{C}^{\prime}}$ and $V_{\mathcal{C}}$ are disjoint, so the union of $\mathcal{C}$ and $\mathcal{C}^{\prime}$ is iterated tripartite, as required. ∎
It remains to prove Lemma 3.2. We first record a technical result about binary trees. To state this result, we define the weight of an internal vertex in a binary tree to be $2$ less than the number of leaves in the subtree consisting of this vertex and all its descendants.
Lemma 3.4.
There exists a constant $C$ such that, for any binary tree $T$ with $k$ leaves and any collection $X$ of vertices of $T$ with total weight less than $k\log k/C$, the number of triples of leaves whose least common ancestor lies outside $X$ is at least $k^{2.5}$.
Proof.
For any vertex $v$, let $n_{v}$ be the number of leaves in the subtree rooted at $v$ and let $m_{v}$ be the number of triples of leaves whose least common ancestor is $v$. Define an imbalance in a tree $T$ to be a configuration where $u$ is the parent of $v$ and $w$, $v$ is the parent of $x$ and $y$, $n_{x}>n_{w}$ and $n_{x}\geq n_{y}$. In this case, we say that the imbalance is at $u$. The rotation of the imbalance is the modified tree $T^{\prime}$ where the subtree rooted at $x$ is swapped with the smaller subtree rooted at $w$. Note that the rotation increases the depths of leaves below $w$ by $1$ and decreases the depths of the leaves below $x$ by $1$. If we set $D_{\textup{total}}$ to be the sum of the depths of the leaves in the binary tree, then the rotation at $u$ decreases $D_{\textup{total}}$ by $n_{x}-n_{w}>0$.
Let $V_{\textup{in}}$ be the set of internal vertices of $T$. For any function $f:V_{\textup{in}}\to[0,1]$ that maps the internal vertices to real numbers in $[0,1]$, we define its weight by $\sum_{v\in V_{\textup{in}}}f(v)\cdot(n_{v}-2)$ and its score by $\sum_{v\in V_{\textup{in}}}(1-f(v))\cdot m_{v}$. To complete the proof of the lemma, it will suffice to show that there exists a constant $C$ such that any function $f:V_{\textup{in}}\to[0,1]$ with weight less than $k\log k/C$ has score at least $k^{2.5}$.
For any $W\geq 0$, let $\nu^{*}_{T}(W)$ be the minimum score of a function $f:V_{\textup{in}}\to[0,1]$ with weight at most $W$. If $T$ has an imbalance labeled with $u,v,w,x,y$ as in the definition, let $T^{\prime}$ be the rotation of the imbalance. Then the only changes in weights or scores occur at $u$ and $v$. The following table records how the weights and scores change at these two vertices:
$T$ (before rotation) $T^{\prime}$ (after rotation)
$n_{u}$ $n_{u}=(n_{x}+n_{y}+n_{w})-2$ $n_{u}^{\prime}=(n_{x}+n_{y}+n_{w})-2=n_{u}$
$n_{v}$ $n_{v}=(n_{x}+n_{y})-2$ $n_{v}^{\prime}=(n_{y}+n_{w})-2$
$m_{u}$ $m_{u}=\frac{1}{2}(n_{x}+n_{y})n_{w}[(n_{x}+n_{y}+n_{w})-2]$ $m_{u}^{\prime}=\frac{1}{2}n_{x}(n_{y}+n_{w})[(n_{x}+n_{y}+n_{w})-2]$
$m_{v}$ $m_{v}=\frac{1}{2}n_{x}n_{y}[(n_{x}+n_{y})-2]$ $m_{v}^{\prime}=\frac{1}{2}n_{y}n_{w}[(n_{y}+n_{w})-2]$
Let $f$ be the optimizer for $\nu^{*}_{T}(W)$. Let $W^{\prime}=f(u)(n_{u}-2)+f(v)(n_{v}-2)\in[0,n_{u}+n_{v}-4]$ be the total contribution of $u$ and $v$ to the weight of $W$. By the optimality of $f$, we know that $(f(u),f(v))$ is the maximizer of the following linear program:
maximize $m_{u}a+m_{v}b$
subject to $\begin{cases}0\leq a,b\leq 1,\\ (n_{u}-2)a+(n_{v}-2)b\leq W^{\prime}.\end{cases}$
Denote by $M(W^{\prime})$ the optimum of this linear program. If we set
[…]
then $M(W^{\prime})$ is a line through the origin with slope $m_{t}/(n_{t}-2)$ on $0,n_{t}-2]$ and the slope becomes $m_{t^{\prime}}/(n_{t^{\prime}}-2)$ on $n_{t}-2,n_{u}+n_{v}-4]$.
We may similarly define $M^{\prime}(W^{\prime})$, where we maximize $m_{u}^{\prime}a+m_{v}^{\prime}b$ subject to the conditions $a,b\in[0,1]$ and $(n_{u}^{\prime}-2)a+(n_{v}^{\prime}-2)b\leq W^{\prime}$. Note now that
$\frac{m_{u}^{\prime}}{n_{u}^{\prime}-2}=\frac{1}{2}(n_{y}+n_{w})n_{x}>\frac{1}{2}n_{x}n_{y}>\frac{1}{2}n_{y}n_{w}=\frac{m_{v}^{\prime}}{n_{v}^{\prime}-2},$
so $M^{\prime}(W^{\prime})$ is a line through the origin with slope $m_{u}^{\prime}/(n_{u}^{\prime}-2)$ on $0,n_{u}^{\prime}-2]$ and the slope becomes $m_{v}^{\prime}/(n_{v}^{\prime}-2)$ on $n^{\prime}_{u}-2,n_{u}^{\prime}+n_{v}^{\prime}-4]$, after which it remains constant.
Given the definitions of the functions $M(W^{\prime})$ and $M^{\prime}(W^{\prime})$, we make the following claim.
Claim 1.
For all $W^{\prime}\in[0,n_{u}+n_{v}-4]$, $M^{\prime}(W^{\prime})\geq M(W^{\prime})$.
Proof.
As $M^{\prime}(W^{\prime})$ is concave and $M(W^{\prime})$ is piecewise linear, it suffices to verify this for $W^{\prime}=0$, $W^{\prime}=n_{t}-2$ and $W^{\prime}=n_{u}+n_{v}-4$. The inequality clearly holds at $W^{\prime}=0$. Moreover, since $n_{t}-2\leq n_{u}-2=n_{u}^{\prime}-2$, we know that
$M^{\prime}\left(n_{t}-2\right)-M(n_{t}-2)=\frac{m_{u}^{\prime}}{n_{u}^{\prime}-2}(n_{t}-2)-\frac{m_{t}}{n_{t}-2}(n_{t}-2)=\left(\frac{m_{u}^{\prime}}{n_{u}^{\prime}-2}-\frac{m_{t}}{n_{t}-2}\right)(n_{t}-2).$
Hence, to verify the inequality for $W^{\prime}=n_{t}-2$, it suffices to show that $m_{u}^{\prime}/(n_{u}^{\prime}-2)\geq m_{t}/(n_{t}-2)$. Computing directly, we see that
$\frac{m_{u}^{\prime}}{n_{u}^{\prime}-2}=\frac{1}{2}(n_{y}+n_{w})n_{x}>\frac{1}{2}(n_{x}+n_{y})n_{w}=\frac{m_{u}}{n_{u}-2}$
and
$\frac{m_{u}^{\prime}}{n_{u}^{\prime}-2}=\frac{1}{2}(n_{y}+n_{w})n_{x}>\frac{1}{2}n_{x}n_{y}=\frac{m_{v}}{n_{v}-2}.$
Finally, to verify the inequality for $W^{\prime}=n_{u}+n_{v}-4$, note that $M(n_{u}+n_{v}-4)$, $M^{\prime}(n_{u}^{\prime}+n_{v}^{\prime}-4)$ are equal as they both count the triples of leaves which meet at least two of the subtrees rooted at $x$, $y$ and $w$. Since $n_{u}^{\prime}+n_{v}^{\prime}-4<n_{u}+n_{v}-4$, we have $M^{\prime}(n_{u}+n_{v}-4)=M^{\prime}(n_{u}^{\prime}+n_{v}^{\prime}-4)=M(n_{u}+n_{v}-4)$, as desired. ∎
If we now set $f^{\prime}$ to be the same as $f$ except that $f^{\prime}(u)=a$, $f^{\prime}(v)=b$, where $(a,b)$ is the extremizer for $M^{\prime}(W^{\prime})$, then both the weight and the score of $f^{\prime}$ in $T^{\prime}$ are at most those of $f$ in $T$, so that $\nu_{T^{\prime}}^{*}(W)\leq\nu_{T}^{*}(W)$. That is, rotating imbalances does not increase $\nu_{T}^{*}(W)$, so, since we are trying to give a lower bound on $\nu_{T}^{*}(W)$, we may rotate as often as we please.
To apply this observation, we need to show that for any binary tree $T$ it is possible to do a series of rotations so that the resulting binary tree has no remaining imbalances. But, as observed earlier, any rotation strictly decreases $D_{\textup{total}}$, the sum of the depths of the leaves. Since $D_{\textup{total}}$ is always a non-negative integer, this shows that the binary tree will have no imbalances after finitely many rotations.
We may therefore assume that $T$ has no imbalances. In particular, for any vertex $u$ with children $v$ and $w$, we must have that $n_{v},n_{w}$ are both in $n_{u}/3,2n_{u}/3]$. Otherwise, without loss of generality, we may assume that $n_{w}<n_{u}/3$ and $n_{v}>2n_{u}/3$. But then there would exist a child $x$ of $v$ such that $n_{x}>n_{u}/3>n_{w}$, contradicting that there are no imbalances. Therefore, any vertex $u$ at depth $d<0.1\log k$ is an internal vertex with weight $\Omega(3^{-d}k)$ and $m_{u}=\Omega(3^{-3d}k^{3})$. Since $\Omega(3^{-3d}k^{3})=\omega(k^{2.5})$, if the score of a function $f:V_{\textup{in}}\to[0,1]$ is less than $k^{2.5}$, then any vertex whose depth is less than $0.1\log k$ must be mapped to a number which is at least $1/2$ by $f$. Since the vertices at depth $d$ have total weight $k-O(2^{d})$ for every $d<0.1\log k$, it follows that the weight of $f$ is $\Omega(k\log k)$, as desired. ∎
Now we are ready for the proof of the key lemma.
Proof of Lemma 3.2..
Let $N=2^{\ell}$ and let $G$ be a complete graph on vertex set $[N]$. For each $i\in[\ell]$ and $u\in[N]$, let $c_{i}(u)$ be an independent uniform element of $[A]$. Finally, for any $u\neq v\in[N]$, if $t=v_{2}(u-v)$, where $v_{2}(x)$ is the $2$-adic valuation function whose value is the highest power of $2$ dividing $x$, we color the edge $uv$ with the product color $(t,(-1)^{\lfloor u/2^{t}\rfloor}(c_{t}(u)-c_{t}(v))\bmod A)$. Note that as $\lfloor u/2^{t}\rfloor$ and $\lfloor v/2^{t}\rfloor$ have different parities, the color of $uv$ does not depend on how we order $u$ and $v$. Moreover, each color class is the union of vertex-disjoint bicliques, as required.
Let $k$ be a positive integer to be determined later and let $S\subseteq[N]$ be any subset of order $k$. We wish to upper bound the probability that there are too few rainbow triangles on this set of vertices. To this end, consider the following binary tree $T$ induced by $S$: place $S$ at the root vertex; consider the first bit where some elements of $S$ do not agree when written in binary and then split $S=S_{0}\cup S_{1}$ based on the value of this bit, where we assume $\left\lvert S_{0}\right\rvert\geq\left\lvert S_{1}\right\rvert$; attach the binary trees induced by $S_{0}$ and $S_{1}$ to the root vertex.
For an internal vertex corresponding to the set $S^{\prime}$, let $S_{0}^{\prime}$ and $S_{1}^{\prime}$ be the two subsets corresponding to its children with $\left\lvert S_{0}^{\prime}\right\rvert\geq\left\lvert S_{1}^{\prime}\right\rvert$. Suppose that the first bit where some elements disagree is the $t$-th bit. Call $S^{\prime}$ good if there is no color that appears more than $|S^{\prime}_{0}|/2$ times among the collection of $c_{t}(u)$ with $u\in S_{0}^{\prime}$ and bad otherwise. If $\left\lvert S^{\prime}\right\rvert\geq 3$, then $\left\lvert S_{0}^{\prime}\right\rvert\geq 2$, so the probability that $S^{\prime}$ is bad can be bounded above by
$A\cdot 2^{|S^{\prime}_{0}|}\cdot A^{-(|S^{\prime}_{0}|+1)/2}\leq\exp(-\Omega((\left\lvert S^{\prime}\right\rvert-2))\log A),$
where we used that $A\geq 20$. Moreover, the event that $S^{\prime}$ is bad is independent from all other similar events.
If $S^{\prime}$ is good, then we know that, for every $u\in S_{0}^{\prime}$, there are at least $\left\lvert S_{0}^{\prime}\right\rvert/2$ vertices $v\in S_{0}^{\prime}$ with $c_{t}(u)\neq c_{t}(v)$. For any such $u,v\in S_{0}^{\prime}$ and any $w\in S_{1}^{\prime}$, it is clear that uvwuvw is rainbow. Therefore, we have found at least
[…]
rainbow triangles with $S^{\prime}$ as their least common ancestor. Note that this is exactly a quarter of the number of triplets whose least common ancestor is $S^{\prime}$.
If we now apply Lemma 3.4 with $X$ the set of bad vertices in $T$, we see that if the total weight on the vertices in $X$ is less than $k\log k/C$, then the number of triples of leaves whose least common ancestor is outside $X$, and therefore good, is at least $k^{2.5}$. By the observation above, this would mean that we have at least $k^{2.5}/4$ rainbow triangles in $S$. We may therefore assume that the total weight on the set of bad vertices is at least $k\log k/C$. Since the weight of the vertex corresponding to $S^{\prime}$ is $\left\lvert S^{\prime}\right\rvert-2$ and the probability that $S^{\prime}$ is bad can be bounded by $\exp(-\Omega((\left\lvert S^{\prime}\right\rvert-2))\log A)$, the probability that all of the vertices in $X$ are bad is at most $\exp(-\Omega(k\log k\log A))$. Since there are at most $2^{k}$ possible choices for $X$ (as there are exactly k1k-1 internal vertices), a union bound implies that the probability there are fewer than $k^{2.5}/4$ rainbow triangles in $S$ is at most $\exp(-\Omega(k\log k\log A))$.
To ensure, by a union bound, that there are at least $k^{2.5}/4$ rainbow triangles in all vertex subsets $S$ of order $k$ with positive probability, we need that
$\binom{N}{k}\exp(-\Omega(k\log k\log A))<1,$
or, equivalently, that $k\log k\log A=\Omega(k\log N).$ Therefore, it suffices to take $k=\exp(\Omega(\log N/\log A))$. ∎
4 Concluding remarks
The main problem left open by this paper is whether Conjecture 1.1 holds in full generality. However, it would already be interesting to prove it for hypergraphs with three tight components or to find a different proof for the two component case that gives a better bound. A particular case of interest is the Fano plane $F$, which is the unique $3$-graph with seven edges on seven vertices in which every pair of vertices is contained in a unique edge. The Fano plane is not iterated tripartite, so, according to Conjecture 1.1, $r(F,K_{n}^{(3)})$ should not grow polynomially. A proof of this would considerably strengthen our belief in the conjecture, on whose validity the authors do not form a completely united front. Let us also point out that in trying to prove 1.1, one cannot hope to prove a statement analogous to 2.1 and 3.1 that gives a single Ramsey construction avoiding the entire family of possible $H$ simultaneously. Indeed, if a red/blue edge coloring of $K_{N}^{(3)}$ has all its red subgraphs iterated tripartite, it is not hard to show by induction that there is a blue clique of size $2^{\log_{3}N+O(1)}=\Omega(N^{\log_{3}2}).$
Another problem of interest is to determine the growth rate of $r(\mathcal{H},K_{n}^{(3)})$, where $\mathcal{H}$ is the family of all non-tripartite tightly connected $3$-graphs. 2.1 gives the lower bound $r(\mathcal{H},K_{n}^{(3)})\geq 2^{\Omega(n^{2/3})}$. As an upper bound, we can show that $r(\mathcal{H},K_{n}^{(3)})\leq 2^{O(n)}$. To see this, for each positive integer $N$, let $n(N)$ be the largest positive integer such that every red/blue-colored $K_{N}^{(3)}$ with no red copy of any hypergraph from $\mathcal{H}$ contains a blue $K_{n(N)}^{(3)}$. We claim that $n(N)\geq 1+n\left(\lceil\frac{N-1}{2}\rceil\right)$. Assuming this claim, a simple induction implies that $n(2^{k}-1)\geq k$ for all integers $k\geq 1$ and we are done.
[…]
While we suspect that the correct bound for $r(\mathcal{H},K_{n}^{(3)})$ is of the form $2^{\Theta(n)}$, it would be more interesting if it were $2^{O(n^{1-\epsilon})}$ for some $\epsilon>0$. Such intermediate growth has been demonstrated [7] for the Ramsey numbers $r(K_{4}^{(3)},S_{n}^{(3)})$, where $S_{n}^{(3)}$ is the $3$-graph on $n+1$ vertices consisting of all $\binom{n}{2}$ edges incident to a given vertex, but it would be very interesting to have such an example with $K_{n}^{(3)}$ instead of $S_{n}^{(3)}$. A particular family of cases where such intermediate growth has not been ruled out is $r(C_{\ell}\setminus e,K_{n}^{(3)})$, where $C_{\ell}\setminus e$ is the tight cycle of length $\ell\not\equiv 0\pmod{3}$ with a single edge removed. While Theorem 1.2 gives $r(C_{\ell}\setminus e,K_{n}^{(3)})\geq 2^{\Omega(n^{2/3})}$, we do not currently know of any upper bound better than $r(C_{\ell}\setminus e,K_{n}^{(3)})\leq r(C_{\ell},K_{n}^{(3)})\leq 2^{O_{\ell}(n\log n)}$ for $\ell$ sufficiently large [14].
It might be interesting to look at the analogous conjecture to Conjecture 1.1 for higher uniformities. If Conjecture 1.1 is indeed correct, its generalization should say that for a $k$-graph $H$, there exists a constant $c$ depending only on $H$ such that $r(H,K_{n}^{(k)})\leq n^{c}$ for all $n$ if and only if $H$ is contained in an iterated blowup of an edge. One direction of this conjecture is simple, while certain partial results in the other direction again hold. For example, a straightforward extension of either construction in Section 2 implies that the conjecture holds for $2$-tightly connected hypergraphs, hypergraphs where any two edges $e$ and $f$ are joined by a sequence of edges $e=e_{0},e_{1},\dots,e_{t}=f$ such that $e_{i-1}$ and $e_{i}$ share at least two vertices for all $i=1,\dots,t$. In particular, if we generalize the argument that yields Theorem 1.2, we obtain the following result.
Theorem 4.1.
If $k\geq 3$ and $H$ is a $k$-graph which is 2-tightly connected and not $k$-partite, then $r(H,K_{n}^{(k)})\geq 2^{\Omega(n^{2/k})}$.
One might also try to find necessary and sufficient conditions on $k$-graphs $H$ under which $r(H,K_{n}^{(k)})$ is upper bounded by a function of the form $2^{n^{c}}$ or $2^{2^{n^{c}}}$ and so on. In this direction, we make the following, likely difficult, conjecture, which would extend Theorem 4.1. Generalizing the definition above, we say that a hypergraph is $s$-tightly connected if any two edges $e$ and $f$ are joined by a sequence of edges $e=e_{0},e_{1},\dots,e_{t}=f$ such that $e_{i-1}$ and $e_{i}$ share at least $s$ vertices for all $i=1,\dots,t$.
Conjecture 4.2.
If $k>s$ and $H$ is a $k$-graph which is $s$-tightly connected and not $k$-partite, then there exists a positive constant $c$ such that $r(H,K_{n}^{(k)})\geq t_{s}(n^{c})$, where the tower function $t_{i}(x)$ is defined by $t_{1}(x)=x$ and $t_{i}(x)=2^{t_{i-1}(x)}$ for all $i\geq 2$.
One natural way to prove 4.2 would be to consider the growth of the Ramsey number $r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$, where $\mathcal{H}_{s}^{(k)}$ is the family of $s$-tightly connected $k$-graphs that are not $k$-partite. We can make an even stronger conjecture than 4.2.
Conjecture 4.3.
If $k>s$, then there exists a positive constant $c$ such that $r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})\geq t_{s}(n^{c})$.
Our methods prove this conjecture for $s=2$ and 4.2 follows immediately from 4.3 as $r(H,K_{n}^{(k)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ for any $H\in\mathcal{H}_{s}^{(k)}$. We note that all the proofs in this paper followed this strategy. Moreover, we observe that 4.3, if true in general, would also determine the tower height of diagonal hypergraph Ramsey numbers, perhaps the central open problem in hypergraph Ramsey theory. Indeed, $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ for any positive integers $s,k,n$ with $k>s$, so we would have that the tower height of $r(K_{n}^{(s)},K_{n}^{(s)})$ is at least s1s-1, agreeing with the longstanding upper bound.
To see the inequality $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$, suppose for the sake of contradiction that there is a red/blue-coloring of the edges of $K_{N}^{(k)}$ with $N=r(K_{n}^{(s)},K_{n}^{(s)})$ that avoids any red $H$ in $\mathcal{H}_{s}^{(k)}$. Label the $k$ parts of each of the red $s$-tight components from $1$ to $k$. Consider the auxiliary coloring of the edges of $K_{N}^{(s)}$ where an edge $e$ is red if it is contained in a red edge $\tilde{e}$ in the original coloring and $e$ contains precisely the vertices in $\tilde{e}$ from parts $1,2,\ldots,s$ in the red $s$-tight component of $\tilde{e}$; $e$ is blue otherwise. Since $\left\lvert e\right\rvert=s$, any red $\tilde{e}$ containing $e$ belongs to the same red $s$-tight component, showing that the color of $e$ does not depend on the choice of $\tilde{e}$. It is clear that each red $\tilde{e}$ in $K_{N}^{(k)}$ contains exactly one red $e$ in the auxiliary coloring of $K_{N}^{(s)}$. By the definition of $r(K_{n}^{(s)},K_{n}^{(s)})$, this auxiliary coloring contains either a red $K_{n}^{(s)}$ or a blue $K_{n}^{(s)}$. We will show that either one would give a blue $K_{n}^{(k)}$ in the original coloring. Indeed, suppose that $S$ is the set of $n$ vertices that forms a monochromatic $K_{n}^{(s)}$ in the auxiliary coloring. Then each $\tilde{e}\in\binom{S}{k}$ contains either 0 or $\binom{k}{s}>1$ red edges $e$, so it has to be blue. This yields the required contradiction and so $r(K_{n}^{(s)},K_{n}^{(s)})\geq r(\mathcal{H}_{s}^{(k)},K_{n}^{(k)})$ must hold.
Acknowledgments.
We are grateful to Jiaxi Nie, Maya Sankar and Yuval Wigderson for stimulating conversations. We are also grateful to the anonymous reviewers for several helpful remarks.